# Image-Based Roadmaps for Vision-Only Planning and Control of Robotic Manipulators

### v1

@@ IV-D Adaptive Visual Servoing
This work builds on the adaptive visual servoing method described in [15], using a roadmap of collision-free sequence of goal image states. At each goal, vector of keypoints ($\mathds{k}$) in image state ($\mathbb{K}$), is tracked as visual features as described in [16]. The controller moves the arm minimizing the feature error, computed as the difference between the current and the target $\mathbb{K}$. The Jacobian matrix, estimated online via least-square optimization of recent joint velocities and keypoints vector over a moving window, eliminates the need to read joint position from encoder. This makes the control pipeline completely model-free. To improve accuracy, we reset the Jacobian estimation window at each new target keeping the estimate unbiased and relevant to the current goal. Since the goals may not be evenly spaced in the image for different experiments, we use a saturation limit on the feature error to prevent sudden spikes in velocity ensuring stable motion and protecting the motor from damage.
Figure 5: Histograms of actual joint displacements along the edges for the three roadmaps. The Learned roadmap closely aligns with the distribution of the Joint Space roadmap, showing only slight deviations and indicating reasonable accuracy in the predicted joint displacements. In contrast, the Image space roadmap demonstrates a wider spread and larger joint displacements, reflecting its lack of alignment between image-space proximity and joint-space movements. This misalignment may reduce its efficiency in generating paths suitable for precise control.
@@ V Experiments and Observations:
We experimentally assesed the performance of the proposed vision only motion planning framework on a Franka Emika Panda Arm.
@@ V-A1 Generating samples
We generated the required image state samples by collecting a large dataset created by actuating the planar joints (Joints $2,4,6$) of the Franka arm, using velocities computed from Eq. 5 as explained in subsection IV-A, covering the robot’s planar workspace.
V-A2 Generating the roadmap
The collected samples $\mathbb{K}$ were used as nodes to generate roadmaps using the different proposed distances described in subsection IV-B. The roadmap generated by the image-space distance is coined as Image Space roadmap and the one generated using the learned distance is coined as Learned roadmap. For benchmarking purposes we also use the actual joint distance to create the Joint Space roadmap. The Joint Space roadmap used actual joint displacements from encoders solely for benchmarking, maintaining the proposed model-free framework. The k-neighbor value for all approaches was set to $25$.
@@ V-A3 Obstacle Representation and Path Planning
In the experimental setup, obstacles were modeled as virtual yellow rectangles with a safety margin, as shown in Fig. 4. Paths were generated offline for various start ($\mathbb{K}_{\text{start}}$) and goal ($\mathbb{K}_{\text{goal}}$) incorporating the obstacle avoidance logic from subsection IV-C. These paths were later used in adaptive visual servoing experiments
V-A4 Control Experiment Set-up
@@ V-B1 Distribution of Joint Distances of Edges for Different Roadmaps
Fig. 5 shows the histograms of joint displacements along the edges of the three roadmaps. The Learned roadmap closely aligns with the joint space roadmap with minor deviations, demonstrating reasonable accuracy in estimated joint displacements. In contrast, the Image Space roadmap exhibits a broader spread and larger joint displacements. This suggests that the Learned roadmap is more effective in accurately capturing transitions and generating paths that minimize joint movements.
@@ V-B2 Average Joint Distances for Planned Collision-Free Paths
We randomly selected $100$ pairs of $\mathbb{K}_{\text{start}}$ and $\mathbb{K}_{\text{goal}}$ from the roadmaps for path planning without obstacles. The average joint distances over $100$ trials, summarized in subsubsection V-B2, show that the learned roadmap yields results closer to the joint space while the image space roadmap results in significantly larger joint displacements. These findings suggest that the learned roadmap possibly generates more efficient path, better suited for successful control convergence.
TABLE I: Average Joint Distances (Radians) for Collision-Free Paths
Joint Space Learned Image
Space
Mean
(radians) 1.74 2.19 3.06
@@ V-B3 Planned paths for Control Experiments
We generated planned paths for two scenarios: $16$ start and goal pairs without checking for collision and $10$ pairs with collision avoidance. Paths were computed using the three roadmaps: Joint Space, Learned, and Image Space. These precomputed paths were used in the control experiments described in subsection V-C.
@@
As observed in Table II the learned roadmap consistently resulted in fewer waypoints and shorter joint distances compared to the image space roadmap, which prioritizes minimizing keypoints distances in image space but incurs higher joint displacements.
@@
TABLE II: Comparison of Joint Distances and Keypoints Distance in Image Space over Experiments
Without Collision-check With Collision-check
@@
  • The larger number of waypoints in the image space roadmap arises from its Euclidean distance-based edge weights, causing A* to prioritize shorter pixel distances and select more intermediate nodes. In contrast, the learned roadmap, with joint displacement-based weights, produces more direct paths with fewer waypoints.
  • The image states $\mathbb{K}$ which are close in image space may be much further away in joint space as highlighted in Table II. This behavior may lead to increased overall joint movement where non-holonomy may exist in the joint space for the covered joint displacements. This may reduce control accuracy and hinder convergence to the target image state.
@@ V-C Control Experiments
The adaptive visual servoing experiments [^2] Planning and Control Experiments videos are available at this link. The details of how to use the link is in the supplementary ReadMe file used precomputed paths from subsubsection V-B3, with control gains optimized to minimize rise and settling times while keeping overshoot within 5%, by careful tuning.
In subsection V-C, the overall control metrics highlight that the image space roadmap succeeded in only $69.2$% of cases, while the learned roadmap achieved a $100$% success rate. However, the image space roadmap, when successful, showed faster transient responses compared to the Learned roadmap. Table IV uses identical metrics as in Table II, categorizing experiments into successful and failed cases
Failures in the image space roadmap were characterized by large joint displacements relative to smaller image space distances, as noted earlier in subsubsection V-B3. Optimal execution of the references generated by using image space distances requires all the keypoints to follow straight paths. This is not possible for the robot due to the non-holonomic constraints of its kinematics (if defined directly in image space). As a result, the robot deviates from the planned path. While the robot reaches the reference locations in most cases, the deviations cause it to go out of its workspace (due to joint limits) in some others. This effect is especially visible when obstacles exist in the workspace since the robot needs to travel near the edges of its workspace to avoid them. When estimated joint distances are used the resulting trajectories are more suitable to the robot’s kinematics, which prevents such failures at large.
TABLE III: Performance results for the control experiments with and without obstacle
Without Collision-check With Collision-check
Performance
Metrics Joint
Space Learned Image
Space Joint
Space Learned Image
Space
Successful
Experiments 16/16 16/16 13/16 10/10 10/10 5/10
System
Rise time (s) 74.92 97.53 94.75 105.38 125.38 90.46
System
Settling time (s) 94.37 118.62 101.99 118.18 155.14 96.44
End effector
Rise time (s) 74.86 97.53 94.39 105.02 125.38 90.46
End effector
Settling time (s) 94.37 114.02 99.81 118.18 147.14 94.22
@@
Execution
time (s) 125.17 148.92 140.78 146.02 201.68 124.70
TABLE IV: Comparison of Joint Distances and Keypoints Distance in Image Space over Experiments for Successful and Failed Experiments
@@
To summarize, both the image space and learned roadmaps exhibit unique benefits. The image space roadmap provides faster execution when successful but struggles in reliability and path convergence. The learned roadmap, using joint displacement-based distances, avoids non-holonomic constraints and ensures robustness, particularly in complex paths. The choice between the two ultimately depends on the specific application context.
@@ VI Conclusion and Future Work
In conclusion, this work introduced a novel framework for collision-free motion planning of robotic manipulators that relied solely on visual features, eliminating the need for explicit robot models or encoder feedback.
The learned roadmap offered smoother, more reliable transitions, and due to its joint displacement-based distance definition, the paths it generated maintained joint-space holonomy when it existed. In contrast, the paths produced by the image space roadmap sometimes failed to maintain joint-space holonomy, even when holonomy existed in the image space. However, the image space roadmap provided faster transient responses and simplicity, making it advantageous for applications where speed and computational efficiency were prioritized.
Future work will explore extending this approach to out-of-plane motion and incorporating real-world obstacles to develop a fully integrated control and manipulation pipeline.

### v2

@@ IV-D Adaptive Visual Servoing
This work builds on the adaptive visual servoing method described in [15], using a roadmap of collision-free sequence of goal image states. At each goal, vector of keypoints ($\mathds{k}$) in image state ($\mathbb{K}$) is tracked as visual features as described in [16]. The controller moves the arm minimizing the feature error, computed as the difference between the current and the target $\mathbb{K}$. The Jacobian matrix, estimated online via least-square optimization of recent joint velocities and keypoints vector over a moving window, eliminates the need to read joint position from encoder. This makes the control pipeline completely model-free. To improve accuracy, we reset the Jacobian estimation window at each new target keeping the estimate unbiased and relevant to the current goal. Since goal spacing in the image may vary across experiments, a saturation limit prevents velocity spikes, ensuring smooth motion and protecting the motor.
Figure 5: Distribution of joint displacements along roadmap edges. The Learned roadmap closely aligns with the Joint Space distribution, indicating accurate distance estimation. In contrast, the Image Space roadmap shows a wider spread, reflecting poor alignment between image and joint space.
@@ V Experiments and Observations:
We experimentally assessed the performance of the proposed vision only motion planning framework on a Franka Emika Panda Arm [37].
@@ V-A1 Generating Samples
We generated the required image state samples by collecting a finite dataset created by actuating the planar joints (Joints $2,4,6$) of the Franka arm, using velocities computed from Eq. 5 as explained in Section IV-A, covering the robot’s planar workspace.
V-A2 Building the Roadmap
The collected samples $\mathbb{K}$ were used as nodes to build roadmaps based on the distance metrics described in Section IV-B. We refer to the roadmap using image space distances as the Image Space roadmap, and the one using learned distances as the Learned roadmap. For benchmarking, we also construct a Joint Space roadmap using actual joint displacements from encoders, used only for comparison, while preserving the model-free design. The k-neighbor value for all approaches was set to $25$. The first row of Table I shows the time required to generate each roadmap. The learned roadmap takes significantly longer to construct due to the additional computational overhead of running neural network inference during nearest neighbor computation.
@@ V-A3 Obstacle Representation and Path Planning
In the experimental setup, obstacles were modeled as virtual shapes such as a rectangle or a triangle or a circle with a safety margin, as shown in the first image of Fig. 4. Paths were generated offline for various start ($\mathbb{K}_{\text{start}}$) and goal ($\mathbb{K}_{\text{goal}}$) incorporating the obstacle avoidance logic from Section IV-C. These paths were later used in adaptive visual servoing experiments.
V-A4 Control Experiment Set-up for Single Obstacle
@@
V-A5 Control Experiment Set-up for Multiple Obstacles
To further evaluate the robustness of our planning method, we conducted $4$ additional experiments involving scenes with multiple obstacles of varying shapes and sizes. The learned roadmap successfully generated feasible paths in all $4$ scenarios. In contrast, the image space roadmap was able to find valid paths in only $2$ out of $4$ cases, failing in the remaining two due to limitations in its Euclidean distance metric when navigating tighter environments.
V-A6 Control Experiment Set-up for Virtual Shelf
We conducted a proof-of-concept experiment simulating a real-world object (a multi-level shelf), as shown in the second image of Fig. 4, and planned paths using both the learned and image space roadmaps.
@@ V-B1 Distribution of Joint Distances of Edges for Different Roadmaps
Fig. 5 shows the joint displacement histograms for the three roadmaps. The learned roadmap closely matches the joint space roadmap, with only minor deviations, indicating accurate estimation of joint displacements. In contrast, the image space roadmap shows a wider spread and larger displacements, suggesting less efficient transitions.
@@ V-B2 Average Joint Distances for Planned Collision-Free Paths
We randomly selected $100$ $\mathbb{K}_{\text{start}}$ and $\mathbb{K}_{\text{goal}}$ pairs from the roadmaps for path planning without obstacles. As shown in Table II, the learned roadmap produces joint distances closer to the joint space baseline, while the image space roadmap results in significantly higher displacements. This indicates that the learned roadmap generates more efficient paths, better suited for control convergence.
TABLE I: Roadmap building and average path generation time (s)
Joint Space Learned Image Space
Roadmap creation time (s) 0.25 766.19 0.28
Query path time (s) $2.422$ $3.19$ $2.38$
TABLE II: Average joint distances (radians) for collision-free paths
Joint Space Learned Image Space
Mean (radians) 1.74 2.19 3.06
@@ V-B3 Planned paths for Control Experiments
We generated planned paths for two scenarios: $16$ start and goal pairs without checking for collision and $10$ pairs with collision avoidance. Paths were computed using the three roadmaps: joint space, learned, and image space. These precomputed paths were used in the control experiments described in Section V-C. The second row of Table I reports the average time each roadmap took to generate paths for these scenarios.
@@
As observed in Table III the learned roadmap consistently resulted in fewer waypoints and shorter joint distances compared to the image space roadmap, which prioritizes minimizing keypoints distances in image space but incurs higher joint displacements.
@@
TABLE III: Comparison of joint distances and keypoints distances in image space over experiments
Without Obstacle Avoidance With Obstacle Avoidance
@@
  • The image space roadmap uses Euclidean distances, leading A* to favor shorter pixel paths with more intermediate waypoints. In contrast, the learned roadmap, based on joint displacements, yields more direct paths with fewer waypoints.
  • Image states that are close in pixel space can be far apart in joint space, as shown in Table III. This can result in larger joint movements and control errors, especially in systems with non-holonomic constraints.
@@ V-C Control Experiments
The adaptive visual servoing experiments [^2] Planning and control experiments including apf videos, are available at this link. The details of how to use the link is in the supplementary ReadMe file used precomputed paths from Section V-B3, with control gains optimized to minimize rise and settling times while keeping overshoot within 5%, by careful tuning. We also performed adaptive visual-servoing experiments for the $4$ scenarios for learned and image space roadmaps as defined in Section V-A5. To highlight the advantages of planning-based control over reactive methods, we implemented an Artificial Potential Field (APF) controller [38] and tested it on the same $4$ multi-obstacle scenarios from Section V-A5.
The control metric in Table IV shows that the image space roadmap succeeded in $69.2$% of cases, while the learned roadmap achieved $100$% success. When successful, the image space roadmap had faster transients. This can be attributed to its shorter pixel-distance paths, which reduce image space error more quickly in scenarios where joint displacements are relatively smaller. Table V uses the same metrics as Table III, grouped by success and failure.
Image space roadmaps fail when the robot tries to follow short pixel paths that require large joint movements. Due to non-holonomic constraints, the robot might need to traverse a curved path in joint space to achieve even a small shift in image space, risking collisions or workspace violations. Learned roadmaps avoid this by producing motions that better match the robot’s capabilities.
Table VI depicts the performance of learned and image space roadmaps in multi-obstacle scenarios. While image space failed to plan in $50$% of the cases, all successful plans led to successful control; the learned roadmap succeeded in both planning and control for all cases. The APF controller failed to converge in all scenarios.
Despite its simplicity, APF failed in all cases, where the robot started between closely spaced obstacles, getting trapped in local minima and oscillating between attractive and opposing repulsive forces, without reaching the goal. These failures highlight the limitations of reactive methods in cluttered scenes and the necessity of global planners like our roadmap-based approach for collision avoidance.
Table VII presents control performance in a proof-of-concept scenario involving simulated real-world object, as described in Section V-A6. Here, the path planned by the image space roadmap led to a collision with the object, despite a defined safety margin. Further experiments are needed to conclusively determine why the robot collided despite a collision-free planned path.
TABLE IV: Performance results for the control experiments with and without obstacle avoidance
Without Obstacle Avoidance With Obstcale Avoidance
Performance Metrics Joint Space Learned Image Space Joint Space Learned Image Space
Successful Experiments 16/16 16/16 13/16 10/10 10/10 5/10
System Rise time (s) 74.92 97.53 94.75 105.38 125.38 90.46
System Settling time (s) 94.37 118.62 101.99 118.18 155.14 96.44
End effector Rise time (s) 74.86 97.53 94.39 105.02 125.38 90.46
End effector Settling time (s) 94.37 114.02 99.81 118.18 147.14 94.22
@@
Execution time (s) 125.17 148.92 140.78 146.02 201.68 124.70
TABLE V: Comparison of joint distances and keypoints distances in image space over experiments for successful and failed experiments
@@
TABLE VI: Performance comparison of learned and image space roadmaps and APF with multiple obstacles
Roadmap Successful Planning. Successful Control. System Rise Time (s) System Settling Time (s) EE Rise Time (s) EE Settling Time (s) Overshoot (%) Exec. Time (s)
Learned 4/4 4/4 96.53 99.68 95.98 97.83 0.87 121.65
Image Space 2/4 2/2 89.1 102.6 89.1 100.05 0.43 122.45
APF NA 0 NA NA NA NA NA NA
TABLE VII: Performance comparison for learned and image space roadmaps with proof-of-concept real-object scenario
Roadmap Successful Planning. Successful Control. System Rise Time (s) System Settling Time (s) EE Rise Time (s) EE Settling Time (s) Overshoot (%) Exec. Time (s)
Learned 1/1 1/1 34.8 75.8 28.9 75.8 3.24 97.5
Image Space 1/1 0/1 NA NA NA NA NA NA
In summary, both roadmaps exhibit unique benifits. The image space roadmap enables faster execution when successful but lacks reliability, while the learned roadmap offers greater robustness by leveraging joint displacement-like distances. The choice depends on the application context.
@@ VI Conclusion and Future Work
In conclusion, this work introduced a novel framework for collision-free motion planning of robotic manipulators that relied solely on visual features, eliminating the need for explicit robot models or encoder feedback. The learned roadmap offered smoother, more reliable transitions, and due to its joint displacement-based distance definition, the paths it generated maintained joint-space holonomy when it existed. In contrast, the paths produced by the image space roadmap sometimes failed to maintain joint-space holonomy, even when holonomy existed in the image space. However, the image space roadmap provided faster transient responses and simplicity, making it advantageous for applications where speed and computational efficiency were prioritized.
Future work will explore extending this approach to out-of-plane motion, incorporating real-world obstacles, and handling occlusions to develop a fully integrated control and manipulation pipeline.